# Thermometer: Towards Universal Calibration for Large Language Models

### 20 Feb 2024 v1

@@ 4.2 Test Time Procedure
$\mathbb{E}\left|\frac{1}{N_{\ast}}\sum_{n=1}^{N_{\ast}}\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}))-\mathbb{E}_{{\bm{x}}\sim\mathcal{P}_{\ast}}\bigg{[}\psi_{\bm{\theta}}(\phi({\bm{x}}))\bigg{]}\right|\leq\frac{4}{3}C_{\theta}\frac{\log N_{\ast}}{N_{\ast}}+(2V_{\theta})^{\frac{1}{2}}\sqrt{\frac{\log N_{\ast}}{N_{\ast}}},\;\;\text{and,}$,
$\mathrm{CE}\left(\frac{1}{N_{\ast}}\sum_{n=1}^{n}\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}))\right)-\mathrm{CE}\bigg{(}\mathbb{E}_{{\bm{x}}\sim\mathcal{P}_{\ast}}\bigg{[}\psi_{\bm{\theta}}(\phi({\bm{x}}))\bigg{]}\bigg{)}\leq L\left(\frac{4}{3}C_{\theta}\frac{\log N_{\ast}}{N_{\ast}}+(2V_{\theta})^{\frac{1}{2}}\sqrt{\frac{\log N_{\ast}}{N_{\ast}}}\right).$
@@ 4.3 Question Answering with Free Form Answers
As we introduced in Section 3, given a sequence of tokens as prompts ${\bm{x}}_{n}$, we aim to calibrate the LLM’s prediction on its corresponding completion token $y_{n}$. In the context of a multiple-choice QA problem, $y_{n}$ typically represents a single token corresponding to one of the multiple-choice options, such as “A" or “B". However, in practice, LLMs are more commonly tasked with generating free-form responses for arbitrary questions without predefined answer options.
@@ 5 Experiments
Methods ECE TL-ECE MCE NLL Brier
[…]
Benchmark Datasets We employ two widely used benchmark datasets for multiple-choice question-and-answer (QA) experiments: MMLU (Hendrycks et al., 2021) and BIG-bench (Srivastava et al., 2022), and adopt MRQA (Fisch et al., 2019) for experiments on QA with free form answers. MMLU contains datasets of exam questions from fifty seven subjects spanning various fields, each question comprises a question context and four possible answers. BIG-bench is a massive benchmark containing more than two hundred datasets covering a wide range of NLP tasks. We extracted multiple-choice QA datasets with over thousand training instances, which resulted in twenty three datasets. The MRQA benchmark comprises of a predefined train and development split, with six training datasets and six development datasets. The style of questions vary greatly both within and across the training and development sets. More details are in Section C.2.
Models We aim to enhance the calibration performance of instruction tuned language models. We consider both encoder-decoder and decoder-only models. For our experiments, we employ FLAN-T5-XL with three billion parameters (Chung et al., 2022) and LLaMA-2-Chat 7B with seven billion parameters (Touvron et al., 2023) as representative examples of encoder-decoder models and decoder-only models. thermometer is an auxiliary model with a Multi-Layer Perceptron (MLP)-like structure described in Section C.3.
Evaluation Metrics We evaluate the calibration performance using following metrics: (1) Expected Calibration Error (ECE): measures the average error between prediction confidence and accuracy across different confidence intervals, for which we use 10 bins in our evaluation. (2) Maximum Calibration Error (MCE) (Naeini et al., 2015): identifies the largest error between prediction confidence and accuracy among all confidence bins, represents the worst-case calibration scenario. (3) Negative log likelihood (NLL): a proper scoring rule (Gneiting and Raftery, 2007) that measures closeness to the data generating process in the Kullback-Leibler sense (Deshpande et al., 2024). (4) Brier Score (Brier) (Brier, 1950): another proper scoring rule that measures the mean squared error between prediction confidence and one-hot labels. (5) Top Label ECE (TL-ECE) (Gupta and Ramdas, 2021): measures the average error conditioned on both confidence and top-predicted labels. TL-ECE also serves as an upper bound of ECE, and is an alternate measure of calibration error.
Baselines Our goal of calibrating models on tasks with no labeled data rules out many calibration approaches. Nonetheless, we compare thermometer against several well-established methods and sensible baselines: (1) Vanilla: calibration performance of LLMs without any calibration techniques applied. (2) Temperature Scaling (TS): cheats by using labeled data from the testing task to tune the task-specific temperatures and establishes a lower bound for the error. (3) Temperature Scaling with cross-validation (TS-CV): This variant of TS is similar to our approach. Here, the temperature is tuned using data from $K-1$ tasks, and used to calibrate the held-out task, without any task-specific adaptation. (4) Monte Carlo Dropout (MC-Dropout(Gal and Ghahramani, 2016): performs inference with random dropout multiple times, and averages the output probabilities across these inferences. (5) Monte Carlo Augmentation (MC-Augment(Wei and Zou, 2019): employs random data augmentations of the prompts rather than dropout. Importantly, we found dropout to not be functional in LLaMA-2-Chat 7B, enabling dropout at inference time did not produced any variability, so we only report MC-Dropout comparisons with FLAN-T5-XL.
5.1 Main Results
Refer to caption Refer to caption
Methods ECE TL-ECE MCE NLL Brier
[…]
Multiple-choice QA
Our main calibration results for multiple-choice QA tasks are presented in two formats. We offer an overview of the overall calibration performance by calculating average results across all datasets. This includes 57 tasks from MMLU and 23 from BIG-bench. The average results and $95\%$ confidence intervals for both FLAN-T5-XL and LLaMA-2-Chat 7B models on the MMLU dataset are in Table 1 and Table 2. We provide a finer granularity view through scatter plots illustrating performance on each task in Figure 9 and Figure 10 in the appendix. In these plots, points above the diagonal are instances where the calibration method fails to improve over the uncalibrated model. Notably, thermometer shows lower number of failures compared to other baselines, with zero failure cases on FLAN-T5-XL. Moreover, compared to the closest competitor, TS-CV, thermometer with LLaMA-2-Chat 7B  produces lower ECE and TL-ECE scores for $50/57$ ($48/57$ with FLAN-T5-XL) MMLU datasets. Furthermore, we compare the predicted temperatures of thermometer, which uses no labeled data from the test task, with temperatures learned via temperature scaling that uses all the labeled data from the task in Figure 2. We find thermometer’s predictions adapt to different tasks and align with learned temperatures.
The diverse nature of the BIG-bench tasks makes learning thermometer more challenging. Nonetheless, the results for BIG-bench, included in Section A.2, show that thermometer with FLAN-T5-XL outperforms all competing approaches, producing lower ECE and TL-ECE scores than the strongest competitor, TS-CV, on $20/23$ tasks. For LLaMA-2-Chat 7B, its performance is within noise of TS-CV, but even then outperforms TS-CV on a majority ($12/23$) of the tasks.
QA with Free Form Answers
We use the MRQA shared task to evaluate thermometer’s efficacy for calibrating QA tasks with free form answers. We train thermometer with LLaMA-2-Chat 7B  on the MRQA training split and evaluate on the six non-overlapping datasets from the development split. The results are presented in Table 3. Similar to MMLU, we find that thermometer substantially improves calibration of the uncalibrated model and performs favorably against competing methods. Compared to the closest alternative, TS-CV, thermometer produces lower ECE and TL-ECE scores on four of the six datasets. These results suggest that thermometer can be effective at calibrating broader (non-QA) free-form generation tasks.
5.2 Analysis
thermometer transfers across model scales and benchmarks
Figure 3: thermometer transfers across model scales. We use FLan-T5-XL 3B thermometer predicted temperatures to calibrate FLan-T5-XL 3B (bold axes), FLan-T5-XXL 11B and Flan-UL2 20B in the top subplots, and LLaMA-2-Chat 7B thermometer predicted temperatures for calibrating LLaMA-2-Chat 7B (bold axes), LLaMA-2-Chat 13B and LLaMA-2-Chat 70B in the bottom subplots. In these plots, each dot represents a MMLU task. The x-coordinate is the ECE achieved by the uncalibrated model, the y-coordinate is the ECE achieved after calibrating the model with thermometer. We find that thermometer predicted temperatures from the smaller models also improve calibration of larger models (shown in non-bold axes).
Figure 5: Inference Runtime Comparison. thermometer is significantly faster than methods that require multiple forward passes and nearly at par with the uncalibrated model.
thermometer incurs low inference overhead
Inference runtime is an important factor for calibrating LLMs. In Figure 5, we compare the inference runtime of different calibration methods on the global facts dataset from MMLU [^6] Similar trends hold across all datasets, which comprises ninety nine data points. Since thermometer only requires an additional forward pass through a MLP, its inference runtime is nearly at par with the uncalibrated LLM and substantially faster than MC-Augment. Although we only report the numbers for MC-Augment, we expect any approach that is asymptotically more accurate with the number of forward passes (for example, MC-Dropout or other approximate Bayesian approaches), to fare similarly. Moreover, thermometer is substantially better at calibrating than the competitors, it achieves an ECE of $0.080\pm 0.008$, while the unclaibrated model’s ECE is $0.260\pm 0.025$, TS-CV is at $0.119\pm 0.014$, and MC-Augment’s ECE is $0.156\pm 0.015$.
thermometer is effective in small labeled data regimes
thermometer demonstrates a significant advantage in calibrating new test tasks by predicting the desired temperature using only unlabeled data. In contrast, TS requires labeled data for temperature tuning. The TS results in Table 2, which utilize all available labeled data, serve as a lower-bound benchmark for our method. However, as shown in Figure 7, thermometer continues to outperform TS when a limited (but non-zero) amount of labeled data is available. This finding suggests that when labeled data for a task is scarce or nonexistent, it is more effective to rely on thermometer for calibration, rather than employing temperature scaling.
thermometer’s performance improves with number of training tasks
Figure 8: Aggregation v.s. Sample-wise Temperature. Sample-wise temperature scaling is less effective compared to thermometer, demonstrating the superiority of the temperature aggregation approach.
6 Concluding Remarks
This work introduces thermometer, a simple yet effective approach for calibrating LLMs. Through comprehensive empirical evaluations we find that it improves calibration, is robust to data shifts, and is computationally efficient. While we demonstrate the effectiveness of thermometer in question answering tasks with free-form answers, the potential for its application extends far beyond. Future directions include adapting thermometer for other complex free-form generation tasks, such as summarization and translation, and applying thermometer to even larger LLMs.
7 Acknowledgements
We are grateful to Natalia Martinez Gil, Dan Gutfreund, Farzaneh Mirzazadeh, and Veronika Thost for their feedback on the manuscript and to Inkit Padhi for help with compute infrastructure.
Appendix A Additional Results
A.1 Additional Results on MMLU
We include additional scatter plots of MMLU datasets in Figure 9 for FLAN-T5-XL and Figure 10 for LLaMA-2-Chat 7B.
@@ A.1 Additional Results on MMLU
Figure 10: LLaMA-2-Chat 7B Scatter Plots: ECE Score of 57 MMLU Datasets. The x-axis and y-axis represent ECE score of uncalibared and calibrated model, respectively. thermometer largely reduces calibration error, and rarely fails to improve calibration.
A.2 Results on BIG-bench
@@
Figure 13: thermometer against temperature scaling on BIG-bench. Comparison of thermometer predicted temperature and optimal temperature obtained by temperature scaling. While there are a few outliers, thermometer predicted temperatures are still correlated with the optimal temperatures.
The average calibration performance of thermometer on the BIG-bench dataset is presented in Table 4 for FLAN-T5-XL and Table 5 for LLaMA-2-Chat 7B. Additionally, scatter plots illustrating these results are detailed in Figure 11 for FLAN-T5-XL and Figure 12 for LLaMA-2-Chat 7B. The comparison between thermometer’s predicted temperatures and the optimal temperatures obtained through temperature scaling is depicted in Figure 13. Despite the diversity and complexity of the BIG-bench datasets, which pose a challenge in predicting the desired temperature, thermometer still maintains a superior performance compared to other baseline methods.
A.3 Ablation Studies

### 27 Jun 2024 v2

@@ 4.2 Test Time Procedure
$\mathbb{E}\left|\frac{1}{N_{\ast}}\sum_{n=1}^{N_{\ast}}\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}))-\mathbb{E}_{{\bm{x}}\sim\mathcal{P}_{\ast}}\bigg{[}\psi_{\bm{\theta}}(\phi({\bm{x}}))\bigg{]}\right|$
$\qquad\leq\frac{4}{3}C_{\theta}\frac{\log N_{\ast}}{N_{\ast}}+(2V_{\theta})^{\frac{1}{2}}\sqrt{\frac{\log N_{\ast}}{N_{\ast}}},\;\;\text{and,}$,
$\mathrm{CE}\left(\frac{1}{N_{\ast}}\sum_{n=1}^{n}\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}))\right)-\mathrm{CE}\bigg{(}\mathbb{E}_{{\bm{x}}\sim\mathcal{P}_{\ast}}\bigg{[}\psi_{\bm{\theta}}(\phi({\bm{x}}))\bigg{]}\bigg{)}$
$\qquad\leq L\left(\frac{4}{3}C_{\theta}\frac{\log N_{\ast}}{N_{\ast}}+(2V_{\theta})^{\frac{1}{2}}\sqrt{\frac{\log N_{\ast}}{N_{\ast}}}\right).$
@@ 4.3 Question Answering with Free Form Answers
As we introduced in Section 3, given a sequence of tokens as prompts ${\bm{x}}_{n}$, we aim to calibrate the LLM’s prediction on its corresponding completion token $y_{n}$. In the context of a multiple-choice QA problem, $y_{n}$ typically represents a single token corresponding to one of the multiple-choice options, such as ‘A’ or ‘B’. However, in practice, LLMs are more commonly tasked with generating free-form responses for arbitrary questions without predefined answer options.
@@
Methods ECE TL-ECE MCE NLL Brier
[…]
Elicitation 0.315$\pm$0.016 0.340$\pm$0.014 0.787 $\pm$0.034 /// ///
Elicitation-Ensemble 0.202$\pm$0.019 0.255$\pm$0.017 0.544$\pm$0.055 /// ///
CAPE 0.196$\pm$0.021 0.247$\pm$0.016 0.431$\pm$0.054 1.542$\pm$ 0.097 0.184$\pm$0.009
Figure 3: LLaMA-2-Chat 7B Scatter Plots: ECE Score of 57 MMLU Datasets. The x-axis and y-axis represent ECE score of uncalibared and calibrated model, respectively. thermometer reduces calibration error on all 57 datasets, often substantially.
@@ 5 Experiments
Benchmark Datasets We employ two widely used benchmark datasets for multiple-choice question-and-answer (QA) experiments: MMLU (Hendrycks et al., 2020) and BIG-bench (Srivastava et al., 2022), and adopt MRQA (Fisch et al., 2019) for experiments on QA with free form answers. MMLU contains datasets of exam questions from fifty seven subjects spanning various fields, each question comprises a question context and four possible answers. BIG-bench is a massive benchmark containing more than two hundred datasets covering a wide range of NLP tasks. We extracted multiple-choice QA datasets with over thousand training instances, which resulted in twenty three datasets. The MRQA benchmark comprises of a predefined train and development split, with six training datasets and six development datasets. The style of questions vary greatly both within and across the training and development sets. More details are in Section C.2.
Models We aim to enhance the calibration performance of instruction tuned language models. As decoder-only models become more ubiquitous in practice, our main experiments are conducted using decoder-only model LLaMA-2-Chat 7B with seven billion parameters (Touvron et al., 2023). To demonstrate the robust calibration performance of our proposed thermometer with respect to different types of models, we also conduct a set of experiments using encoder-decoder model FLAN-T5-XL with three billion parameters (Chung et al., 2022). thermometer is an auxiliary model with a Multi-Layer Perceptron (MLP)-like structure described in Section C.3.
Evaluation Metrics We evaluate the calibration performance using following metrics: (1) Expected Calibration Error (ECE): measures the average error between prediction confidence and accuracy across different confidence intervals, for which we use 10 bins in our evaluation. (2) Maximum Calibration Error (MCE) (Naeini et al., 2015): identifies the largest error between prediction confidence and accuracy among all confidence bins, represents the worst-case calibration scenario. (3) Negative log likelihood (NLL): a proper scoring rule (Gneiting & Raftery, 2007) that measures closeness to the data generating process in the Kullback-Leibler sense (Deshpande et al., 2024). (4) Brier Score (Brier) (Brier, 1950): another proper scoring rule that measures the mean squared error between prediction confidence and one-hot labels. (5) Top Label ECE (TL-ECE) (Gupta & Ramdas, 2021): measures the average error conditioned on both confidence and top-predicted labels. TL-ECE also serves as an upper bound of ECE, and is an alternate measure of calibration error.
Baselines Our goal of calibrating models on tasks with no labeled data rules out many calibration approaches. Nonetheless, we compare thermometer against several classical calibration methods as well as those recently proposed methods: (1) Vanilla: calibration performance of LLMs without any calibration techniques applied. (2) Temperature Scaling (TS): cheats by using labeled data from the testing task to tune the task-specific temperatures and establishes a lower bound for the error. (3) Temperature Scaling with cross-validation (TS-CV): This variant of TS is similar to the setting of our approach. Here, the temperature is tuned using data from $K-1$ tasks, and used to calibrate the held-out task, without any task-specific adaptation. (4) Monte Carlo Dropout (MC-Dropout) [^8] We found dropout to not be functional in LLaMA-2-Chat 7B, enabling dropout at inference time did not produced any variability, so we only report MC-Dropout comparisons with FLAN-T5-XL. (Gal & Ghahramani, 2016): performs inference with random dropout multiple times, and averages the output probabilities across these inferences. (5) Monte Carlo Augmentation (MC-Augment(Wei & Zou, 2019): employs random data augmentations of the prompts rather than dropout. (6) Elicitation (Tian et al., 2023) directly generates verbalized confidence by providing the LLMs with carefully crafted prompts. (7) Elicitation-Ensemble (Xiong et al., 2023) improves the quality of verbalized confidence by aggregation of multiple samples. (8) CAPE (Jiang et al., 2023a) performs prompt augmentation using option permutation.
5.1 Main Results
Methods ECE TL-ECE MCE NLL Brier
[…]
Elicitation 0.130$\pm$0.031 0.207$\pm$0.036 0.522$\pm$0.169 /// ///
Elicitation-Ensemble 0.171$\pm$0.047 0.237$\pm$0.032 0.552$\pm$0.206 /// ///
CAPE 0.067$\pm$0.016 0.098$\pm$0.028 0.156$\pm$0.060 0.556$\pm$ 0.087 0.183$\pm$0.036
Figure 5: thermometer transfers across different model scales of LLaMA-2-Chat. We use LLaMA-2-Chat 7B thermometer predicted temperatures for calibrating LLaMA-2-Chat 7B (bold axes), LLaMA-2-Chat 13B and LLaMA-2-Chat 70B. In these plots, each dot represents a MMLU task. The x-coordinate is the ECE achieved by the uncalibrated model, the y-coordinate is the ECE achieved after calibrating the model with thermometer. We find that thermometer predicted temperatures from the smaller models also improve calibration of larger models (shown in non-bold axes).
Multiple-choice QA
Our main calibration results for multiple-choice QA tasks are presented in two formats. We offer an overview of the overall calibration performance by calculating average results across all datasets. This includes 57 tasks from MMLU and 23 from BIG-bench. The average results and $95\%$ confidence intervals for MMLU are presented in Table 1. We provide a finer granularity view through scatter plots illustrating performance on each task in Figure 3. In this plots, points above the diagonal are instances where the calibration method fails to improve over the uncalibrated model, where thermometer shows zero failure cases. Furthermore, we compare the predicted temperatures of thermometer, which uses no labeled data from the test task, with temperatures learned via temperature scaling that uses all the labeled data from the task in Figure 4. We find thermometer’s predictions adapt to different tasks and align with learned temperatures.
The diverse nature of the BIG-bench tasks makes learning thermometer more challenging. Nonetheless, the results for BIG-bench, included in Section A.1, thermometer still outperforms other baselines with large margin, while stays within noise of TS-CV. Finally, we remark that the similar behavior of thermometer holds for encoder-decoder model FLAN-T5-XL (see Appendix A.2). thermometer shows superior performance over other baseline methods. For the challenging BIG-bench, thermometer with FLAN-T5-XL outperforms all competing approaches, producing lower ECE and TL-ECE scores than the strongest competitor, TS-CV, on $20/23$ tasks.
QA with Free Form Answers
We use the MRQA shared task to evaluate thermometer’s efficacy for calibrating QA tasks with free form answers. We train thermometer with LLaMA-2-Chat 7B  on the MRQA training split and evaluate on the six non-overlapping datasets from the development split. The results are presented in Table 2. Similar to multiple-choice QA tasks, we find that thermometer substantially improves calibration of the uncalibrated model and performs favorably against competing methods. Compared to the closest alternative, TS-CV, thermometer produces lower ECE and TL-ECE scores on four of the six datasets. These results suggest broader applicability of thermometer to (non-QA) free-form generation tasks.
5.2 Analysis
thermometer transfers across model scales and benchmarks
thermometer is effective in small labeled data regimes
thermometer demonstrates a significant advantage in calibrating new test tasks by predicting the desired temperature using only unlabeled data. In contrast, TS requires labeled data for temperature tuning. The TS results in Table 1, which utilize all available labeled data, serve as a lower-bound benchmark for our method. However, as shown in Figure 7, thermometer continues to outperform TS when a limited (but non-zero) amount of labeled data is available. This suggests that when labeled data for a task is scarce or nonexistent, it is more effective to rely on thermometer than employing temperature scaling.
thermometer’s performance improves with number of training tasks
Figure 9: Aggregation v.s. Sample-wise Temperature. Sample-wise temperature scaling is less effective compared to thermometer.
6 Concluding Remarks
This work introduces thermometer, a simple yet effective approach for calibrating LLMs. Through comprehensive empirical evaluations we find that it improves calibration, is robust to data shifts, and is computationally efficient. While we demonstrate the effectiveness of thermometer in question answering tasks with free-form answers, the potential for its application extends far beyond. Future directions include adapting thermometer for other complex free-form generation tasks, such as summarization and translation, and applying thermometer to larger LLMs.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgements
We are grateful to Natalia Martinez Gil, Dan Gutfreund, Farzaneh Mirzazadeh, and Veronika Thost for their feedback on the manuscript and to Inkit Padhi for help with compute infrastructure
Appendix A Additional Results
A.1 thermometer Consistently Performs Well on BIG-bench
The average calibration performance of thermometer on the BIG-bench dataset is presented in Table 3. Additionally, scatter plots illustrating these results are detailed in Figure 10. The comparison between thermometer’s predicted temperatures and the optimal temperatures obtained through temperature scaling is depicted in Figure 11. Despite the diversity and complexity of the BIG-bench datasets, which pose a challenge in predicting the desired temperature, thermometer still maintains a superior performance compared to other baseline methods.
Methods ECE TL-ECE MCE NLL Brier
[…]
Elicitation 0.225$\pm$0.048 0.263$\pm$0.059 0.745$\pm$0.085 /// ///
Elicitation-Ensemble 0.202$\pm$0.049 0.241$\pm$0.063 0.753$\pm$0.068 /// ///
CAPE 0.214$\pm$0.073 0.235$\pm$0.076 0.475$\pm$0.107 1.567$\pm$0.365 0.231$\pm$0.043
Figure 11: thermometer against temperature scaling on BIG-bench (LLaMA-2-Chat 7B). Comparison of thermometer predicted temperature and optimal temperature obtained by temperature scaling. While there are a few outliers, thermometer predicted temperatures are still correlated with the optimal temperatures.
A.2 thermometer Consistently Performs Well on Encoder-decoder LLMs
To demonstrate thermometer’s robustness with respect to different types of language models, we also conduct a set of additional experiments on encoder-decoder model FLAN-T5-XL with three billion parameters. We empirically observe that elicitation-based calibration methods (Tian et al., 2023; Xiong et al., 2023) are not even functional on FLAN-T5-XL model thus omit reporting their results, i.e., by providing the crafted prompts, FLAN-T5-XL often fails to generate the corresponding verbalized confidence. This indicates that elicitation-based methods (Tian et al., 2023; Xiong et al., 2023) that rely on the ability of a model to faithfully follow nuanced instructions generally do not perform well for smaller and mid-sized open source models. In fact, the authors in (Tian et al., 2023) concede that even for Llama2-70b, “The verbal calibration of the open source model Llama-2-70b-chat is generally weaker than that of closed source models”.
We evaluate the calibration performance of different methods on both MMLU and BIG-bench. The average calibration performance of thermometer on MMLU datasets and BIG-bench datasets are shown in Table 4 and Table 5, respectively. We also provide the scatter plots of comparison for MMLU and BIG-bench in Figure 12 and Figure 13, respectively. Finally, The comparison between thermometer’s predicted temperatures and the optimal temperatures obtained through temperature scaling is presented in Figure 14. Similar to decode-only model LLaMA-2-Chat 7B, thermometer consistently outperforms other baselines. Furthermore, we further validate that thermometer also transfers well across different model scales of Flan-T5 in Figure 16.
@@ Appendix A Additional Results
Figure 14: thermometer against temperature scaling on MMLU and BIG-bench (FLAN-T5-XL). Comparison of thermometer predicted temperature and optimal temperature obtained by temperature scaling. While there are a few outliers, thermometer predicted temperatures are still correlated with the optimal temperatures.
Figure 15: thermometer transfers across different model scales of FLan-T5. We use FLan-T5-XL 3B thermometer predicted temperatures to calibrate FLan-T5-XL 3B (bold axes), FLan-T5-XXL 11B and Flan-UL2 20B. In these plots, each dot represents a MMLU task. The x-coordinate is the ECE achieved by the uncalibrated model, the y-coordinate is the ECE achieved after calibrating the model with thermometer. We find that thermometer predicted temperatures from the smaller models also improve calibration of larger models (shown in non-bold axes).
A.3 thermometer Shows Stronger Transfer-ability Than TS-CV
In Figure 5 and Figure 6, we demonstrate the strong transfer-ability of thermometer across different model scales and datasets. This is the another advantage of thermometer over those inference-based methods (Wei & Zou, 2019; Tian et al., 2023; Xiong et al., 2023; Jiang et al., 2023a), which do have the capability to transfer. TS-CV is the only baseline that has transfer-ability, so we also conduct an additional experiment to compare the transfer-ability of thermometer with the TS-CV baseline. As illustrated Figure 16, Thermometer typically shows stronger transfer-ability than TS-CV, producing lower calibration errors.
Figure 16: Compare Transfer-ability of thermometer with TS-CV. Top: We use FLan-T5-XL 3B thermometer predicted temperatures to calibrate 20B Flan-UL2 in the left subplot, and LLaMA-2-Chat 7B thermometer predicted temperatures for calibrating LLaMA-2-Chat 13B in the right subplot. Bottom: Applying LLaMA-2-Chat 7B thermometer trained on BIG-bench calibrates MMLU and vice-versa.
A.4 Ablation Studies